Smith and Sons Inc. are determining the viability of a new product line. The new product will require a $280,000 piece of equipment. Shipping and installation will cost $20,000. The equipment has a 3-year tax life, and the allowed depreciation for such property are 33%, 45%, 15%, and 7% for Years 1 through 4. Inventory will increase by $15,000, account payable increasing by $8,000 and account receivables increasing by $10,000. The product line is expected to generate annual revenue of $106,000 per year, with cost of goods sold being 36,000 per year and other costs (excluding depreciation) of $12,000 per year. The tax rate is 30 percent, annual interest expense is $11,000 per year, and the required return for this project is 12 percent.

a. Find the year 0 cash flow, FCF0

b. Find the year 2 cash flow, FCF2
Answer #1

Cash Flow year 0 = Equipment cost + shipping cost + Increase in net operating working capital

Equipment cost = 280000

Shipping cost = 20000

Increase in net operating working capital = Increase in inventory + Increase in receivables - increase in payables

= 15000 + 10000 - 8000 = 17000

Cash flow year 0 = 280000 + 20000 + 17000 = 317000

Cash flow year 2 = (Revenues - Costs) * (1 - tax rate) + Depreciation * tax rate

= (106000 - 36000) * (1 - 0.30) + (280000+20000) * 45% * 30%

= 89500

Depreciation is calculated at the sum of equipment cost and shipping costs = (280000 + 20000) = 300000